# Landscape of Thoughts: Visualizing the Reasoning Process of Large Language Models

### v1

@@ Abstract
Numerous applications of large language models (LLMs) rely on their ability to perform step-by-step reasoning. However, the reasoning behavior of LLMs remains poorly understood, posing challenges to research, development, and safety. To address this gap, we introduce landscape of thoughts-the first visualization tool for users to inspect the reasoning paths of chain-of-thought and its derivatives on any multi-choice dataset. Specifically, we represent the states in a reasoning path as feature vectors that quantify their distances to all answer choices. These features are then visualized in two-dimensional plots using t-SNE. Qualitative and quantitative analysis with the landscape of thoughts effectively distinguishes between strong and weak models, correct and incorrect answers, as well as different reasoning tasks. It also uncovers undesirable reasoning patterns, such as low consistency and high uncertainty. Additionally, users can adapt our tool to a model that predicts the property they observe. We showcase this advantage by adapting our tool to a lightweight verifier that evaluates the correctness of reasoning paths. The code is publicly available at: https://github.com/tmlr-group/landscape-of-thoughts.
@@ 1 Introduction
Figure 1: Landscape of thoughts for visualizing the reasoning steps of LLMs. The darker regions in landscapes indicate more thoughts, with Refer to caption indicating incorrect answers and Refer to caption marking correct answers. Given a question, we sample a few thoughts from an LLM and divide them into two categories based on correctness. We visualize the landscape of each category by projecting the thoughts into a two-dimensional feature space, where each density map reflects the distribution of states at a reasoning step. With these landscapes, users can easily discover the reasoning patterns of an LLM or a decoding algorithm. In addition, a predictive model is applied to predict the correctness of landscapes and can help improve reasoning.
@@
Large language models (LLMs) have revolutionized the paradigm of solving problems with their broad spectrum of capabilities. In particular, several useful applications of LLMs, such as tool use (Schick et al., 2023), retrieval-augmented generation (Lewis et al., 2020), and agents (Yao et al., 2023b), heavily rely on their capability of step-by-step reasoning (Wei et al., 2022; Kojima et al., 2022). Although many base models, e.g., OpenAI o1 (Jaech et al., 2024), and decoding algorithms, e.g., test-time scaling-up search (Snell et al., 2024), have been introduced to advance the performance of LLMs on these applications, the underlying reasoning behavior of LLMs remains unclear to the community. This not only hinders the development of algorithms but also poses potential risks at deployment (Anwar et al., 2024).
A few attempts (Wang et al., 2023a; Saparov and He, 2023; Saparov et al., 2023; Dziri et al., 2024) have been made to understand the reasoning capacity of LLMs. Nevertheless, these findings are often tied to certain decoding algorithms and tasks, which may not be so instructive for users working with their own algorithms and tasks. Instead, there is a strong demand for tools that can be applied to analyze the reasoning behavior of LLMs in the users’ scenarios. We foresee that such tools will benefit three groups of practitioners: 1) engineers can iterate their solutions faster based on the feedback from the tool; 2) researchers can improve decoding algorithms based on insights revealed by the tool; 3) safety researchers can utilize the tool to monitor, understand, and improve the behavior of LLMs.
@@
We made a small but meaningful step towards the above goal by introducing the landscape of thoughts, a tool for visualizing the reasoning paths produced by chain-of-thought and other step-by-step reasoning algorithms. Given any multi-choice reasoning dataset, our tool visualizes the distribution of intermediate states and any reasoning path of interest w.r.t. the answer choices, which enables users to uncover reasoning patterns of LLMs in both success and failure cases (Fig. 1). The core idea is to characterize the textual states in a reasoning path as features that quantify their distances to all answer choices. These distances are estimated by the perplexity metric, with the same LLM to generate thoughts and explain for itself. The state features are then projected to a two-dimensional space via t-SNE (van der Maaten and Hinton, 2008), a non-linear dimensionality reduction method to preserve manifolds in the original high-dimensional space.
@@
We examine our tool with different combinations of model sizes, decoding algorithms, and benchmark datasets. Our tool reveals several qualitative observations regarding the reasoning behaviors of LLMs. Some notable observations include: 1) the convergence speed of reasoning paths towards correct answers reflects the accuracy, no matter what base model, decoding algorithm, or dataset is used; 2) the convergence speed of reasoning paths in success and failure cases is distinct, indicating that we may use the convergence speed of a reasoning path to predict its accuracy; 3) low consistency and high uncertainty are generally observed in the intermediate thoughts, presenting the unstable properties of the reasoning process. To our knowledge, these observations have not been reported by previous works that analyze chain-of-thought mostly based on performance metrics.
@@
Since our tool is built on the top of state features, it can be adapted to a machine-learning model to quantitatively predict certain properties, such as the findings mentioned above. We showcase this advantage by training a lightweight model to predict the success and failure cases, which is equivalent to verifiers commonly used in LLM reasoning (Cobbe et al., 2021). Even though this verifier is lightweight compared to most LLM-based verifiers, it consistently improves the reasoning performance on most combinations of models, decoding algorithms, and datasets in our experiments. Hence, users can further leverage this advantage to predict other potential properties that they discover in their own scenarios.
@@
  • We introduce the first visualization tool for inspecting the reasoning dynamics of different LLMs and decoding algorithms on any multi-choice reasoning dataset (Sec. 2).
  • Our tool reveals several observations regarding the reasoning behaviors of different models, algorithms, and datasets, offering new insights into the reasoning (Sec. 3).
  • Our tool can also be adapted to a model to predict certain properties and guide the reasoning process, improving LLM reasoning without modifying parameters (Sec. 4).
2 Visualizing Multi-step Reasoning of LLMs
This section outlines a general framework for language models and reasoning algorithms compatible with our tool (Sec. 2.1), demonstrates how it visualizes reasoning by projecting thoughts into a two-dimensional space (Sec. 2.2), and introduces metrics for quantitative analysis (Sec. 2.3).
@@ 2.1 Problem Formulation
Our goal is to visualize the reasoning process of LLMs across a variety of problem types. To achieve this, we aim for a formulation that is sufficiently general to encompass a wide range of use cases. Specifically, we focus on datasets consisting of multiple-choice questions, where each sample $(x,y,{\mathcal{C}})$ comprises a question $x$, a correct answer $y$, and a finite set of candidate choices ${\mathcal{C}}=\{c_{j}\}_{j=1}^{k}$, all represented in textual format. The visualization tool applies to the following language models and reasoning algorithms.
@@
Language models. To explore the landscape of thoughts generated by an LLM $p_{\text{LLM}}(\cdot)$, it is necessary for the model to produce diverse reasoning paths for solving a given problem. This requires the LLM to support sampling during inference $\widehat{y}\sim p_{\text{LLM}}(y|x,{\mathcal{C}})$. For chain-of-thought reasoning, thoughts are sampled autoregressively as $\widehat{t}_{i}\sim p_{\text{LLM}}(t_{i}|x,{\mathcal{C}},\widehat{t}_{1},\ldots,\widehat{t}_{i-1})$. Namely, each thought $\widehat{t}_{i}$ is conditioned on the problem $x$, the candidate set ${\mathcal{C}}$, and the sequence of preceding thoughts ${\widehat{t}_{1},\ldots,\widehat{t}_{i-1}}$. To characterize intermediate states within these reasoning paths, the LLM must also function as a likelihood estimator, enabling the computation of the probability $p_{\text{LLM}}(\widehat{y}|x,{\mathcal{C}},\widehat{t}_{1},\ldots,\widehat{t}_{i})$ of any generation $\widehat{y}$. These two requirements are generally satisfied by most open-source LLMs, such as Llama (Dubey et al., 2024), Mistral (Jiang et al., 2024), and DeepSeek (Liu et al., 2024). However, proprietary LLMs, such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), are excluded as they do not support likelihood estimation.
@@
Reasoning algorithms. While there are many approaches to solving reasoning problems with LLMs (Creswell et al., 2022; Kazemi et al., 2023), this work focuses on chain-of-thought (CoT) (Wei et al., 2022) and its derivatives (Zhou et al., 2023; Yao et al., 2023a), owing to their widespread use and development. These decoding algorithms generally guide the model in generating a structured path of intermediate reasoning thoughts before arriving at the final answer. Note that to visualize a large number of reasoning thoughts effectively, these thoughts should be automatically parsed into distinct units (e.g., via sentence tokenization). This requirement is typically satisfied by most variants of CoT.
2.2 Landscape of Thoughts
@@
Given a collection of reasoning paths generated by an LLM, our tool seeks to visualize how different paths lead to either correct or incorrect answers within a two-dimensional (2D) space, as illustrated in Fig. 1. A key challenge lies in the absence of a direct mapping from the textual space of thoughts to 2D coordinates. To address this gap, we first utilize the same LLM to represent intermediate states as numerical vectors. These state vectors are then projected into a 2D space for visualization. For simplicity, we use the notation $t_{i}$ instead of $\widehat{t}_{i}$, which is clear in the following.
@@
Characterizing the states. Here, the intermediate thoughts $\{t_{i}\}_{i=1}^{n}$ in a reasoning path naturally define a sequence of states $\{s_{i}\}_{i=0}^{n}$, where $s_{0}=[x]$ and $s_{i}=[x,t_{1},t_{2},\ldots,t_{i}]$. Here, we propose to characterize the states as feature vectors using the likelihood function of the LLM. Specifically, the $k$-dim feature vector ${\bm{s}}_{i}$ for state $s_{i}$ is defined as follows:
${\bm{s}}_{i}=[d(s_{i},c_{1}),d(s_{i},c_{2}),\ldots,d(s_{i},c_{k})]^{\top},$ (2.1)
@@
where $d(s_{i},c_{j})$ measures the distance between state $s_{i}$ and choice $c_{j}$. In this context, the vector ${\bm{s}}_{i}$ indicates the relative distances from the state $s_{i}$ to all possible choices $\{c_{j}\}_{j=1}^{k}$. To reduce the effect of length on choices, we implement the distance calculation of $d(s_{i},c_{j})$ through the perplexity metric (Shannon, 1948; Manning, 1999) shown as below:  [^1] The perplexity can also be equivalently expressed as $\text{PPL}(c_{j}|s_{i})=\exp\left(-\frac{1}{|c_{j}|}\sum_{t=1}^{|c_{j}|}\log p_{\text{LLM}}(c_{j}[t]|s_{i},c_{j}[:t])\right)$.
$d(s_{i},c_{j})=p_{\text{LLM}}(c_{j}|s_{i})^{-1/|c_{j}|},$ (2.2)
@@
where $|c_{j}|$ is the number of tokens in $c_{j}$, and $p_{\text{LLM}}(c_{j}|s_{i})$ is the accumulated probability in an autoregressive manner. We further normalize the vector ${\bm{s}}_{i}$ to have a unit L1 normalization. Additionally, to represent the choices as landmarks in the visualization, it is necessary to encode the choices as feature vectors. Notably, we observe that perplexity decreases as the model’s prediction confidence increases. To align with this observation, we define the feature vector ${\bm{c}}_{j}$ for a choice $c_{j}$ in a manner consistent with the perplexity:
${\bm{c}}_{j}=\frac{1}{k}[\text{1}(j\neq 1),\ldots,\text{1}(j\neq k)]^{\top}.$ (2.3)
For $r$ paths, each with $n$ states, we compute the feature vectors for all $r\cdot n$ states.  [^2] Our tool can also be applied to paths with different numbers of states. We assume $n$ states for demonstration purposes. Together with the feature vectors of $k$ choices, we obtain a feature matrix ${\bm{S}}\in{\mathbb{R}}^{k\times(r\cdot n+k)}$
${\bm{S}}=[{\bm{s}}^{(1)}_{1},\ldots,{\bm{s}}^{(1)}_{n},\ldots,{\bm{s}}^{(r)}_{1},\ldots,{\bm{s}}^{(r)}_{n},{\bm{c}}_{1},\ldots,{\bm{c}}_{k}].$ (2.4)
Note that a sufficiently large number of paths is necessary to generate a comprehensive visualization of the reasoning landscape. However, visualizing all samples in a dataset under this setting incurs a significant computational cost. In practice, we found it more efficient to visualize $d$ paths with $\frac{r}{d}$ samples projected into the same space. This approach retains much of the visualization quality while substantially reducing the number of paths required for each sample. The key idea is to rearrange the order of choices such that the correct answer consistently aligns with the same dimension in the $k$-dimensional feature space across all the $r$ samples.
@@
Visualization. After constructing the feature matrix ${\bm{S}}$, we project the states and choices into a 2D space for visualization. This dimensionality reduction step can be accomplished using various existing algorithms (Pearson, 1901; van der Maaten and Hinton, 2008; McInnes et al., 2018). In this study, we employ t-SNE (van der Maaten and Hinton, 2008) due to its ability to preserve the underlying manifolds of the original high-dimensional space and its robustness to a wide range of transformations. By applying t-SNE to the $k$-dim ${\bm{S}}$, we obtain the 2-dim coordinates $\bar{{\bm{S}}}\in{\mathbb{R}}^{2\times(rn+k)}$.
The coordinates of the states define a discrete density function in the 2D space. To create a more intuitive and visually interpretable representation, we smooth this density function using a Parzen window estimator (Silverman, 2018). The smoothed density at a given coordinate $\bar{{\bm{v}}}$ is as follows, where the $\sigma$ controls the radius of Gaussian kernels.
$p(\bar{{\bm{v}}})=\frac{1}{rn}\sum_{\bar{{\bm{s}}}\in\bar{{\bm{S}}}}\exp\left(-\frac{||\bar{{\bm{v}}}-\bar{{\bm{s}}}||^{2}}{2\sigma^{2}}\right).$ (2.5)
2.3 Metrics

### v3

@@ Abstract
Numerous applications of large language models (LLMs) rely on their ability to perform step-by-step reasoning. However, the reasoning behavior of LLMs remains poorly understood, posing challenges to research, development, and safety. To address this gap, we introduce landscape of thoughts (LoT), the first landscape visualization tool to inspect the reasoning trajectories with certain reasoning methods on any multi-choice dataset. We represent the textual states in a trajectory as numerical features that quantify the states’ distances to the answer choices. These features are then visualized in two-dimensional plots using t-SNE. Qualitative and quantitative analysis with the landscape of thoughts effectively distinguishes between strong and weak models, correct and incorrect answers, as well as different reasoning tasks. It also uncovers undesirable reasoning patterns, such as low consistency and high uncertainty. Additionally, users can adapt LoT to a model that predicts the property they observe. We showcase this advantage by adapting LoT to a lightweight verifier that evaluates the correctness of trajectories. Empirically, this verifier boosts the reasoning accuracy and the test-time scaling effect. The code is publicly available at: https://github.com/tmlr-group/landscape-of-thoughts.
@@ 1 Introduction
Figure 1: Landscape of thoughts for visualizing the reasoning steps of LLMs. Note that the red landscape represents wrong reasoning cases, while the blue indicates the correct ones. The darker regions in landscapes indicate more thoughts, with Refer to caption indicating incorrect answers and Refer to caption marking correct answers. Specifically, given a question with multiple choices, we sample a few thoughts from an LLM and divide them into two categories based on correctness. We visualize the landscape of each category by projecting the thoughts into a two-dimensional feature space, where each density map reflects the distribution of states at a reasoning step. With these landscapes, users can easily discover the reasoning patterns of an LLM or a decoding method. In addition, a predictive model is applied to predict the correctness of landscapes and can help improve the accuracy of reasoning.
@@
Large language models (LLMs) have revolutionized the paradigm of solving problems. Many practical applications, e.g., LLM as agent (Schick et al., 2023; Lewis et al., 2020; Yao et al., 2023b), critically depend on step-by-step reasoning (Wei et al., 2022; Kojima et al., 2022). Despite progress in advanced models like OpenAI o1 (Jaech et al., 2024) and decoding methods such as test-time scaling (Snell et al., 2024), the underlying reasoning behavior of LLMs remains poorly understood, hindering the development of these models and posing deployment risks (Anwar et al., 2024).
A few pioneer attempts (Wang et al., 2023a; Saparov & He, 2023; Saparov et al., 2023; Dziri et al., 2024) probe LLM reasoning, but their insights often hinge on specific decoders and tasks. In practice, practitioners debug by manually reading the reasoning trajectories generated by LLMs, which has two drawbacks: (i) scalability—human inspection does not scale (e.g., at 30s per trajectory, 100 trajectories require 50min); and (ii) aggregation—deriving reliable, dataset-level conclusions (e.g., from 10,000 trajectories) is difficult, yielding subjective and even biased summaries. These costs compound during iterative development, where fast, interpretable feedback is essential. Consequently, there is a clear need for general, reusable tools to analyze LLM reasoning in users’ settings. This tool can potentially benefit engineers by speeding iteration, reasoning researchers by informing decoder improvements, and safety researchers by monitoring and improving model behavior.
To this end, we introduce the landscape of thoughts (LoT), a visualization of LLM reasoning trajectories that delivers automatic, objective analysis from single examples to full datasets. Analogous to the t-SNE (van der Maaten & Hinton, 2008), LoT highlights structure in high-dimensional reasoning space. By pairing qualitative landscapes with quantitative metrics (consistency, uncertainty, and perplexity), LoT enables comparison and reveals insights beyond manual inspection.
@@
Specifically, given any multi-choice reasoning dataset, LoT visualizes the distribution of intermediate states in any reasoning trajectories of interest w.r.t. the answer choices, which enables users to uncover reasoning patterns in both success and failure trajectories (Fig. 1). The core idea is to characterize the states of textual thoughts in a trajectory as numerical features that quantify the states’ distances to the answer choices. These distances are estimated by the perplexity metric, with the same LLM to generate thoughts and explain to itself. Then, these state features (i) produce three metric plots and (ii) are projected into a two-dimensional space with t-SNE to generate the landscape.
@@
We examine LoT with different dimensions of model sizes, decoding methods, and reasoning datasets. LoT reveals several insightful observations regarding the reasoning behaviors of LLMs. Some notable observations include: 1) The convergence speed of trajectories towards correct answers reflects the accuracy, no matter what base model, decoding method, or dataset is used; 2) The convergence speed of trajectories in success and failure cases is distinct, indicating that we may use the convergence speed of a reasoning trajectory to predict its accuracy; 3) Low consistency and high uncertainty are generally observed in the intermediate thoughts, presenting the unstable properties of the reasoning process. LoT reveals them by bridging localized text understanding with global reasoning dynamics. To our knowledge, these patterns have not been reported by prior analyses of reasoning, which primarily rely on manual text inspection or aggregate performance metrics.
@@
Since our tool is built on top of state features, it can be adapted to a machine-learning model to quantitatively predict certain properties, such as the findings mentioned above. We showcase this advantage by training a lightweight model to predict the success and failure cases, which is equivalent to verifiers commonly used in LLM reasoning (Cobbe et al., 2021). Even though this verifier is lightweight compared to most LLM-based verifiers, it consistently improves the reasoning performance on most combinations of models, decoding methods, and datasets in our experiments. Hence, users can further leverage this advantage to predict the properties in their scenarios.
@@
  • We introduce the first tool for automatic and scalable visualization of the LLM reasoning procedure, applicable to any open-source models and decoding methods on multi-choice datasets (Sec. 2).
  • Our tool reveals several observations regarding the reasoning behaviors of different language models, decoding methods, and reasoning datasets, offering several new insights (Sec. 3).
  • Our tool can also be adapted to a model to predict certain properties and guide the reasoning process, improving LLM reasoning without modifying the model parameters (Sec. 4).
2 Landscape of Thoughts
@@ 2.1 Problem Formulation
Our goal is to visualize the reasoning trajectories of LLMs across a variety of task domains. Specifically, we target datasets consisting of multiple-choice questions, where each datapoint $(x,y,{\mathcal{C}})$ comprises a question $x$, a correct answer $y$, and a finite set of candidate choices ${\mathcal{C}}=\{c_{j}\}_{j=1}^{k}$, all represented in texts.  [^1] LoT is positioned for multi-choice questions. Appendix A.7 discusses its extension to open-ended tasks. The visualization tool applies to the following models and methods.
@@
Language models. To explore the landscape of thoughts generated by an LLM $p_{\text{LLM}}(\cdot)$, the model should produce diverse reasoning trajectories for solving a problem. In each trajectory, the reasoning thoughts are decoded autoregressively as $\hat{t}_{i}\sim p_{\text{LLM}}(t_{i}|x,{\mathcal{C}},\hat{t}_{1},\ldots,\hat{t}_{i-1})$: each thought $\hat{t}_{i}$ is conditioned on the question $x$, the candidate set $\mathcal{C}$}, and the sequence of preceding thoughts $\hat{t}_{1},\ldots,\hat{t}_{i-1}$}. To characterize intermediate states within these trajectories, the LLM must also function as a likelihood estimator, enabling the computation of the probability $p_{\text{LLM}}(\hat{y}|x,{\mathcal{C}},\hat{t}_{1},\ldots,\hat{t}_{i})$ of any answer $\hat{y}$. These two requirements are generally satisfied by open-source LLMs, such as Llama (Dubey et al., 2024) and DeepSeek (Liu et al., 2024). However, closed-source LLMs like GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023) are excluded, as their likelihood estimation is not publicly supported.
@@
Reasoning methods. While there are many approaches to solving reasoning problems with LLMs (Creswell et al., 2022; Kazemi et al., 2023), this work focuses on chain-of-thought (CoT) (Wei et al., 2022) and its derivatives (Zhou et al., 2023; Yao et al., 2023a), owing to their widespread use and development. These decoding methods generally guide the model in generating a structured trajectory of intermediate reasoning thoughts before arriving at the final answer. To visualize a large number of reasoning thoughts effectively, these thoughts should be automatically parsed into distinct units (e.g., via sentence tokenization). This requirement can be satisfied by most LLMs.  [^2] We empirically verify the robustness of LoT if this requirement does not hold (please see Appendix D.9).
2.2 Qualitative Visualization with Landscapes
@@
Given a collection of reasoning trajectories generated by an LLM, our tool seeks to visualize how different trajectories lead to either correct or incorrect answers within a two-dimensional (2D) space, as illustrated in Fig. 1. A key challenge lies in the absence of a direct mapping from the textual space of thoughts to numerical 2D coordinates. To address this gap, we utilize the same LLM to represent intermediate states as numerical features. These state features are then projected into a 2D space for visualization. For simplicity, we denote a thought as $t_{i}$ instead of $\hat{t}_{i}$, which is clear in the following.
@@
Characterizing the states. Here, the intermediate thoughts $\{t_{i}\}_{i=1}^{n}$ in a reasoning trajectory naturally define a sequence of states $\{s_{i}\}_{i=0}^{n}$, where $s_{0}=[x]$ and $s_{i}=[x,t_{1},t_{2},\ldots,t_{i}]$. Here, we propose to characterize the states as features using the likelihood function of the LLM. Specifically, the $k$-dim feature $\bm{f}_{i}$ for state $s_{i}$ indicates the relative distances from the state $s_{i}$ to all possible choices $\{c_{j}\}_{j=1}^{k}$:
$\bm{f}_{i}\triangleq[d(s_{i},c_{1}),d(s_{i},c_{2}),\ldots,d(s_{i},c_{k})]^{\top},$ (1)
@@
where $d(s_{i},c_{j})$ measures the distance between state $s_{i}$ and choice $c_{j}$. To reduce the effect of length on choices, we calculate $d(s_{i},c_{j})$ through the perplexity metric (Shannon, 1948; Manning, 1999):
$d(s_{i},c_{j})\triangleq\exp\left(-\frac{1}{|c_{j}|}\sum_{t=1}^{|c_{j}|}\log p_{\text{LLM}}(c_{j}[t]|s_{i},c_{j}[:t])\right)=p_{\text{LLM}}(c_{j}|s_{i})^{-1/|c_{j}|},$ (2)
where $|c_{j}|$ is the number of tokens in $c_{j}$, and $p_{\text{LLM}}(c_{j}|s_{i})$ is the accumulated probability in an autoregressive manner. Assume $|c_{j}|=T$, we have $p_{\text{LLM}}(c_{j}|s_{i})=p_{\text{LLM}}(c_{j}[1]|s_{i})\cdot p_{\text{LLM}}(c_{j}[2]|s_{i},c_{j}[1])\cdot p_{\text{LLM}}(c_{j}[3]|s_{i},c_{j}[1],c_{j}[2])\ldots p_{\text{LLM}}(c_{j}[T]|s_{i},c_{j}[1],c_{j}[2]\ldots c_{j}[T-1])$. The token-level probabilities are normalized over the entire vocabulary; $c_{j}[1]$ is the first token of $c_{j}$, and $c_{j}[T]$ is the last token.
@@
We further normalize the vector $\bm{f}_{i}$ to have a unit $\ell_{1}$ normalization. Additionally, to represent the choices as landmarks in the visualization, it is necessary to encode the choices as feature vectors. Notably, the perplexity decreases as the model’s prediction confidence increases. To align with this observation, we define the feature vector $\bm{f}^{c}_{j}$ for a choice $c_{j}$ as:
$\bm{f}^{c}_{j}\triangleq\frac{1}{k}[\text{1}(j\neq 1),\ldots,\text{1}(j\neq k)]^{\top}.$ (3)
For $r$ trajectories, each with $n$ states, we compute the feature vectors for all $r\cdot n$ states. The indicator function $\text{1}(j\neq 1)$ will output 1 if j is not 1, and output 0 if and only if j is 1. Here, the element 0 in $c_{1}$ indicates this choice has 0 distance to the first anchor, which is $c_{1}$ itself. Besides, elements 1 indicate that the distances among anchors are assumed to be the same.  [^3] LoT can be applied to trajectories with different numbers of states. We assume $n$ states for demonstrations. Together with the feature vectors of $k$ choices, we obtain a feature matrix ${\bm{F}}\in{\mathbb{R}}^{k\times(r\cdot n+k)}$ as:
${\bm{F}}\triangleq[\bm{f}^{(1)}_{1},\ldots,\bm{f}^{(1)}_{n},\ldots,\bm{f}^{(r)}_{1},\ldots,\bm{f}^{(r)}_{n},\bm{f}^{c}_{1},\ldots,\bm{f}^{c}_{k}].$ (4)
Note that a sufficiently large number of trajectories is necessary to generate a comprehensive visualization of the reasoning landscape. For computational efficiency, we sample $d$ trajectories per question across all questions, yielding $r=d×|\text{questions}|$ total trajectories. We then normalize feature vectors by reordering choices so the correct answer appears in the first dimension across all questions. As such, we can visualize the landscape of multiple questions by putting their trajectories together, which is more efficient than visualizing by generating enough trajectories per question.
@@
Visualization. After constructing the feature matrix $\bm{F}$}, we project the states and choices into a 2D space for visualization. This step can be accomplished using various existing methods of dimensionality reduction (Pearson, 1901; van der Maaten & Hinton, 2008; McInnes et al., 2018a). We employ t-SNE (van der Maaten & Hinton, 2008) due to its ability to preserve the underlying manifolds of the original high-dimensional space and its robustness to a wide range of transformations.  [^4] Appendix D.8 shows that LoT is compatible and robust with different methods of dimensionality reduction. By applying t-SNE to the $k$-dim $\bm{F}$}, we obtain the 2-dim coordinates $\bar{{\bm{F}}}\in{\mathbb{R}}^{2\times(rn+k)}$. The two dimensions are reduced from the original space, which represents all possible answers, and each state’s coordinates show its distance from different answers. Finally, the coordinates of the states define a discrete density function in the 2D space, presented by the color depth in landscapes.
2.3 Quantitative Visualization with Metrics